Stinson Inc imposes a payback cutoff of 3 years for its international investment projects. Assume the company has the following two projects available:

Year. Cash flow A. Cash flow B

0. -$48,000. -$93000

1. 18,500. 20,500

2. 24,800. 25,500

3. 20,500. 33,500

4. 6,500. 247,000

what is the payback period for each project?

which, if either, projects should the company accept?
Answer #1

The cash flows are:

For project A:
Year 0:-48000
Year 1:18500
Year 2:24800
Year 3:20500
Year 4:6500

Cumulative net cash flows are:
Year 0:-48000
Year 1:18500-48000=-29500
Year 2:24800-29500=-4700
Year 3:20500-4700=15800
Year 4:6500

Payback period=2+4700/20500=2.229 years

For project B:
Year 0:-93000
Year 1:20500-93000=-72500
Year 2:25500-72500=-47000
Year 3:33500-47000=-13500
Year 4:247000-13500=233500
Payback period=3+13500/247000=3.054656 Years

Given that the company imposes a payback cutoff of 3 years, so project with payback of less than 3 years should be accepted

Hence, project A should be accepted